OpenStudy (anonymous):

How would you solve this: V= 4/3*3.14*6?

OpenStudy (anonymous):

=12*3.14*6 =37.68*6 =226.08

OpenStudy (anonymous):

actually it's 1) = (4)/(3) 2) = (1.33)*(3.14) 3) =(4.186)*(6) 4) = 25.12 You always start from left to right while dealing with such problems. Especially multiplication and division, so you start of by dividing 4/3, then mulitplying the quotient with 3.14 and so on and so forth.

OpenStudy (anonymous):

but it if your question was dividing 4 with (3*3.14*6) then the answer is approximately = 0.0708